media%2F66e%2F66e9605c-38af-44d6-82da-ae
media%2F90b%2F90b0e9bf-4613-4926-9c4d-80
media%2F1c7%2F1c7b2ef0-b21e-4838-b71a-e8


Use the regression line to estimate the number of trials it would take to leam these tricks if a dolphin received five treats per trick. Y for X _ 5 would be The head dolphin trainer wants to save money by cutting down on the number of treats the dolphins get. She has asked you to use the least-squares regression line to predict how fast the dolphins can leam tricks if you were to give them zero treats when teaching new tricks. Which of the following is an appropriate response? O The regression line predicts that the dolphins will need O attempts to learn a trick if they are given no treats O Since the regression line predicts that the dolphins will need a negative number of attempts, you can assume the dolphins need 0 attempts if given no treats. O The least-squares regression line was calculated using 1 to 4 treats and should not be used to predict what would happen if the dolphins were given no treats O The regression line predicts that the dolphins will need 12 attempts to leam a trick if they are given no treats Continse without saving
Answer #1

Finding Regression equation -

1 0 946 2067 Y 4547 X3241

Σχ 10, ΣΥ 20, Σχ.Y 45, ΣΧ2 30 Y- 20 r.Y- 45 X2ー30

ΣΥΣΧ2 _ Σχ.ΣΧΥ 20-30-10-45 . 7.5 4.30- 102 n.ΣΧΥΣχ.ΣΥ 4-4510-20 n.ΣΧ2 _ (ΣΧ)2 (10)2 4.30

media%2Fb4e%2Fb4ed5954-7155-4720-9bcc-28

Y hat when x = 5,

Y^ = 7.5 - 1*5 = 7.5-5 = 2.5 ~ 3

These many attempts needed

X Y Y^ Residual
3 4 5 -1
2 5 6 -1
4 4 4 1
1 7 7 1

Answer to MCQ -

Correct option -

a) The regression line predicts that the dolphins will need 0 attempts to learn a trick if they are given no treats.

The graph of the regression line is y-7.5-1 x 8.0 7.5 7.0 até 6.0 5.5 5.0 4.5 4.0 3.5 3.0 2.5 2.0 1.5 -1.0-0.5 0.0 0.5 1.0 1.